# ETF: An Entity Tracing Framework for Hallucination Detection in Code Summaries

### v2

@@ 7.3 Error Analysis
public static int writeShort(ArrayView target, int offset, short value) {...} Summary:  Here’s a more detailed implementation of the above method: ….{JAVA CODE}….. Error-Analysis: Creative Summary by Mistral 7B
@@
Error Case 2- Changed Entity Form: The language models may not refer to the exact names of code entities during generation of the summary. In the example given below, the summary mentions the entity ‘PreparedStatement" as ‘prepared statement" which may not be captured during the named entity recognition phase due to changed entity form. The verification of these kinds of summary may not reflect the inaccuracies due such sentences.
protected PreparedStatement setObjects(Connection connection, String query, Object... objects) throws SQLException {....} Summary:  ……set the parameters of the prepared statement using the provided objects. This method is likely responsible for iterating over the objects and setting them as parameters in the prepared statement. Finally, the method returns the prepared statement object….. Error-Analysis: Changed Entity Form
from Llama3-70B
@@ 8 Conclusion and Future Work
Our work addresses the critical challenge of detecting hallucinations in code summarization, a task that demands a deep understanding of both programming and natural languages. By introducing a novel dataset and the Entity Tracing Framework (ETF) with 73% F1 score, we establish a systematic approach to grounding code entities within summaries, enabling a more interpretable and accurate evaluation of explanations. In future, this framework can be enhanced by incorporating a multi-agent system, leveraging multiple LLMs in tandem to improve prediction accuracy. Additionally, the current framework can be further developed to better mitigate the occurrence of hallucinations.
@@ 9 Limitations
While the framework is intended to be generic, certain components—such as the code parsers for entity detection—are tailored to specific programming languages. Additionally, the performance of large language models (LLMs) can be highly sensitive to the prompts used. Despite significant time and effort spent optimizing prompts for each model, there may still be unexplored prompts that could lead to improved performance. Furthermore, the experimental results presented are based on Java code from the CodeXGLUE benchmark. As a result, additional research is required to verify whether the approach generalizes effectively to other programming languages.

### v3

@@ 7.3 Error Analysis
public static int writeShort(ArrayView target, int offset, short value) {...} Summary:  Here’s a more detailed implementation of the above method: ….{JAVA CODE}….. Creative Summary by Mistral 7B
@@
Error Case 2- Changed Entity Form: Language models may not use the exact names of code entities in summaries. For example, the entity "PreparedStatement" may be referred to as "prepared statement," which could be missed during named entity recognition due to the change in form. The verification of these kinds of summaries may not reflect the inaccuracies due to such sentences.
protected PreparedStatement setObjects(Connection connection, String query, Object... objects) throws SQLException {....} Summary:  ……set the parameters of the prepared statement using the provided objects. This method is likely responsible for iterating over the objects and setting them as parameters in the prepared statement. Finally, the method returns the prepared statement object….. Changed Entity Form by Llama3-70B
@@ 8 Conclusion and Future Work
Our work addresses the critical challenge of detecting hallucinations in code summarization, a task that demands a deep understanding of both programming and natural languages. By introducing a novel dataset and the Entity Tracing Framework (ETF) with 73% F1 score, we establish a systematic approach to grounding code entities within summaries, enabling a more interpretable and accurate evaluation of explanations. In the future, this framework can be enhanced by incorporating a multi-agent system and leveraging multiple LLMs in tandem to improve prediction accuracy. The current framework can be further developed to better mitigate the occurrence of hallucinations.
@@ 9 Limitations
While the framework is designed to be generic, certain components—such as code parsers for entity detection—may be unavailable for low-resource programming languages like COBOL or Perl. Moreover, the performance of large language models (LLMs) heavily depends on their parameter size and the volume of training data. Given the complexity of this task, smaller open-source models may struggle to perform effectively, reinforcing the need for larger LLMs. However, these larger models may not be scalable and often demand significantly greater computational resources.
@@ Appendix C NER Evaluation
Appendix D String Matching Heuristics
By string matching heuristics, we mean character-level matching with the following regex expressions:
  • The word is either preceded by or succeeded by a space char
  • ignore the “” characters since some of the entities are enclosed using these quoted marks by models.
  • Account for brackets: some of the function names in the summary include “()” and some of the variables include “[]”
The above regexes are designed to capture all the cases of entity forms in summary. It can be noted that these regex rules are evaluated in a single module.
Appendix E Annotation Details
We discuss our annotation process and the annotator’s guidelines here:
E.1 Background
The dataset was annotated by eight annotators who are experts in Java and held at least a Master’s degree in Computer Science, with some having a PhD in the field. On average, the annotators had 4+ years of experience in Java programming.
E.2 Guidelines
The annotation process was conducted in two stages. In the first stage, we implemented a three-step procedure to annotate hallucinations in code summaries independent of specific hallucination categories.
The first part of the annotation process involved validating the Named Entity Recognition Output. This involved annotating missed or incorrectly identified entities in the summary itself by selecting the appropriate label from the drop-down.
The second part of the annotation involved evaluating if each sentence accurately describes the entity in the code snippet, marking the entity-sentence pair as:
  • CORRECT: If the relevant sentence correctly describes the code
  • INCORRECT: If the relevant sentence incorrectly describes the code
  • IRRELEVANT: If the relevant sentence does not talk about the mapped entity itself
The third part involved rating the summary based on hallucination severity as
  • POOR (Most part is hallucinated): The generated code summary shows below-average correctness.
  • FAIR (Only some part is hallucinated): The generated code summary meets expectations.
  • GOOD (Almost no hallucination): The generated code summary is completely correct.
The second stage involved defining hallucination categories based on annotator feedback and organizing them into a structured taxonomy (Figure 2). This finalized taxonomy was then provided to the annotators, who were asked to assign a specific hallucination category from the predefined options. Annotators were strongly encouraged to include comments explaining their annotations, as these explanations can be useful for researchers utilizing our dataset.